OpenStudy (anonymous):

jane sells 8 times as many Volvos as Melissa. If the difference in the sales is 35, how many cars did Jane sell?

OpenStudy (anonymous):

make two equations: x = jane y = melissa x = 8y 35 = x + y

OpenStudy (anonymous):

so subsititute the 8y into the second equation for x

OpenStudy (anonymous):

@jazy it should be "35 = x - y" in 2nd eqn, shouldn't it?

OpenStudy (anonymous):

Oh yeah! thanks for correcting me:)

OpenStudy (anonymous):

so we still substitute the value of x into the second equation i think: 35 = (8y) - y 35 = 8y - y 35 = 7y /7 y = 5 now you can find x by putting 5 in for the y-value

OpenStudy (anonymous):

35 = x - 5 Once you figure out what x is you know how many cars jane sold:) I leave it to you!